OpenStudy (anonymous):

Solve the following system of equations: x - 2y = 14 x + 3y = 9 (1, 12) (-1, -12) (12, -1) (12, 1)

OpenStudy (texaschic101):

do you prefer substitution or elimination ?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

C

OpenStudy (anonymous):

Can you work step by step with me? so i can fully understand?

OpenStudy (anonymous):

(12,-1)

OpenStudy (texaschic101):

x - 2y = 14 x + 3y = 9 --->(-1) -------------------- x - 2y = 14 -x - 3y = -9 (result of multiplying by -1) --------------------add -5y = 5 y = -1 now sub -1 in for y in either equation to find x x - 2y = 14 x - 2(-1) = 14 x + 2 = 14 x = 14 - 2 x = 12 answer is : (12,-1) --- sheelo_mughal is correct

OpenStudy (anonymous):

I cant, i just do everything in my head

OpenStudy (anonymous):

why wasnt the 2 added to the 14? thats where i messed up.

OpenStudy (texaschic101):

because when a number goes across the equal sign, it changes its sign....thus, making the 2 become -2

OpenStudy (anonymous):

x+2=14 so when to passes = its sign will automatically change to - x=14-2 x=12

OpenStudy (anonymous):

Okay i didnt want the answer but thank you, the mistake i made was not changing the sign. I kept it the same.

OpenStudy (anonymous):

Just remember the rules of solving equation.

OpenStudy (anonymous):

Thanks. I asked for them to help. If you read what i just said you would know the facts.

OpenStudy (anonymous):

Thanks, when eliminating with 2 variables do i just keep it the same?

OpenStudy (anonymous):

no if u want to move a variable across the = then its sign must change

OpenStudy (anonymous):

+ will become - - will become + X will become\[\div \] \[\div\] will become X

OpenStudy (anonymous):

Thanks can i show you 1 more i just did and can you check to see if what i have is right?

OpenStudy (anonymous):

ya ofcourse

OpenStudy (anonymous):

If you were solving a system of equations and you came to a statement like 1 = 3, what do you know about the solution to the system? The solution is (1, 3) The solution is x = 1 and y = 3 There is no solution There are infinitely many solutions I got C. (No solution)

OpenStudy (anonymous):

is it 1=3 or (1,3)

OpenStudy (anonymous):

just 1=3

OpenStudy (anonymous):

if its y=3 then u can say its B but if its 1=3 then C is right

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

ur welcome